According to the Empirical rule, what percent of the data should be between 174 and 846 for a population with mean of 510 and standard deviation of 168?
Answer by Boreal(15235)

846-510=336
510-174=336
the points are each 2 sd from the mean, since 336/168 = 2
Therefore 95% of the data should be between them assuming normality.